A cone has a base radius of 3x units and a height of 4x units. The surface area of the cone is 1944 times 3.14 square units. Find the value of x. Explain your steps
Answer by rothauserc(4718)

surface area(SA) of a cone = (s * pi *r) + (pi * r^2), where s = slant side, r = radius, pi = 3.14
we are given SA = (1944 * 3.14), r = 3x, h = 4x, then s^2 = (3x)^2 + (4x)^2 and s = 5x, therefore
(1944 * 3.14) = (5x * 3.14 * 3x) + (3.14 * 9x^2)
divide both sides of = by 3.14
1944 = 15x^2 + 9x^2
1944 = 24x^2
81 = x^2
x = 9